OpenStudy (thcsbabygirl1224):

An apple pie has a 12-inch diameter and the angle of one slice is 60 degrees. What is the area of that slice of pie? Use 3.14 for pi . 14.13 square inches 75.36 square inches 3pi square inches 18.84 square inches 0.52 square inches

OpenStudy (thcsbabygirl1224):

@AloneS.

OpenStudy (mathstudent55):

A full circle has how many degrees?

OpenStudy (alones.):

Yes, hold on.

OpenStudy (thcsbabygirl1224):

360

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Correct. The slice is 60 degrees. 60 degrees is what part of 360 degrees?

OpenStudy (alones.):

Since the whole pie has 360 degrees, 60 degrees would be 60/360 which is 1/6 of the pie. The entire pie has an area of 3.14*6^2= 113.04 1/6 of 113.04 is 18.84

OpenStudy (alones.):

Oops, I didn't mean to give out the answer like that.

OpenStudy (mathstudent55):

60/360 = 1/6 Since 60 out of 360 is 1/6, the area of the slice is 1/6 the area of the circle. Use the formula for the area of a circle, and then divide by 6.

OpenStudy (thcsbabygirl1224):

18.84 then if so can you help me on my next too

OpenStudy (alones.):

Sure.

OpenStudy (alones.):

Just be sure to post it in a new post.